# OceanGPT: A Large Language Model for Ocean Science Tasks

### v6

@@ Abstract
Ocean science, which delves into the oceans that are reservoirs of life and biodiversity, is of great significance given that oceans cover over 70% of our planet’s surface. Recently, advances in Large Language Models (LLMs) have transformed the paradigm in science. Despite the success in other domains, current LLMs often fall short in catering to the needs of domain experts like oceanographers, and the potential of LLMs for ocean science is under-explored. The intrinsic reason may be the immense and intricate nature of ocean data as well as the necessity for higher granularity and richness in knowledge. To alleviate these issues, we introduce OceanGPT, the first-ever LLM in the ocean domain, which is expert in various ocean science tasks. We propose DoInstruct, a novel framework to automatically obtain a large volume of ocean domain instruction data, which generates instructions based on multi-agent collaboration. Additionally, we construct the first oceanography benchmark, OceanBench, to evaluate the capabilities of LLMs in the ocean domain. Though comprehensive experiments, OceanGPT not only shows a higher level of knowledge expertise for oceans science tasks but also gains preliminary embodied intelligence capabilities in ocean technology [^1] Codes, data and checkpoints will soon be available at https://github.com/zjunlp/KnowLM..
@@ 1 Introduction
Ocean science, which delves into the intricacies of oceans that cover over 70% of our planet’s surface, is essential not only for understanding the rich reservoirs of life and biodiversity but also for recognizing their pivotal role in regulating the global climate and supporting economies (esaias1998overview; falkowski2012ocean; visbeck2018ocean; DBLP:journals/tgrs/JinHWYGZZP23). Recently, advances in Large Language Models (LLMs) (openai2023gpt4; DBLP:journals/corr/abs-2305-09645; DBLP:journals/corr/abs-2307-08674; DBLP:journals/corr/abs-2306-13549; DBLP:journals/corr/abs-2303-18223) have transformed the paradigm in science domains such as medical science (Moor), molecular science (Fang), protein science (doi:10.1126/science.ade2574) and geoscience (DBLP:journals/corr/abs-2306-05064). However, the potential for the LLM in ocean science is under-explored.
@@
Figure 1: Capabilities of OceanGPT.
Through extensive experiments, OceanGPT shows superiority for diverse ocean science tasks. Note that our benchmark data is based on criteria manually evaluated by ocean experts, and can accurately reflect the capabilities that LLMs possess in the field of ocean science. As depicted in Figure 1, our model can comprehensively answer questions according to the instructions of oceanographers, which demonstrates its expertise in oceanography. We further explore the potential of OceanGPT from the perspectives of ocean engineering. Specifically, we integrate ocean robotics instructions into the training data and evaluate the ability of OceanGPT via code or console commands. OceanGPT not only demonstrates a higher level of knowledge expertise but also gains preliminary embodied intelligence capabilities in ocean technology.
@@ 2 Related Work
The landscape of LLM (DBLP:conf/nips/BrownMRSKDNSSAA20; DBLP:journals/corr/abs-2204-02311; DBLP:journals/corr/abs-2302-13971; DBLP:journals/corr/abs-2307-09288) has rapidly evolved and achieved a series breakthroughs. DBLP:journals/corr/abs-2112-11446; DBLP:journals/corr/abs-2205-01068; DBLP:journals/corr/abs-2201-08239; DBLP:journals/corr/abs-2211-05100; DBLP:conf/iclr/ZengLDWL0YXZXTM23 have explored the performance across a wide range of model scales and broadened the application scope. LLMs also exhibit capabilities in various aspects, including reasoning and knowledge interaction (DBLP:journals/corr/abs-2305-13068; knowlm; DBLP:conf/acl/QiaoO0CYDTHC23; DBLP:journals/corr/abs-2308-11432; DBLP:journals/corr/abs-2309-07864). To obtain LLMs, instruction tuning (DBLP:conf/iclr/WeiBZGYLDDL22; DBLP:journals/corr/abs-2308-10792; DBLP:conf/nips/Ouyang0JAWMZASR22) is a crucial technique to alignment with user preferences and desired outputs. alpaca uses self-instruct (DBLP:conf/acl/WangKMLSKH23) to generate data from manually instructions. vicuna2023; DBLP:journals/corr/abs-2304-12244 uses user-shared conversations and AI-generated data for tuning. Different from those, we introduce an automatic and effective domain instruction generation framework via multi-agent collaboration.
@@
LLMs have emerged as cornerstone models in addressing challenges within scientific research. DBLP:journals/corr/abs-2212-13138 explores the potential of clinical LLMs and introduces a human evaluation framework and instruction prompt tuning. Moor proposes generalist medical AI that is capable of handling diverse medical tasks using self-supervised learning on large datasets. DBLP:journals/corr/abs-2107-03134 introduces MedGPT, a model using EHR data and Named Entity Recognition tools for predicting future medical events. BioGPT (DBLP:journals/bib/LuoSXQZPL22) is a language model pre-trained on biomedical literature for improved text generation and mining. Transfer-biology describes Geneformer, a model pre-trained on single-cell transcriptomes for making predictions with limited data in network biology. doi:10.1126/science.ade2574 demonstrates the prediction of atomic-level protein structure from primary sequences using scaled-up language models. DBLP:journals/corr/abs-2306-05064 introduces the first LLM specifically designed for geoscience, including its training and benchmarking protocols. Different from previous works, we are the first LLM for ocean science tasks and explore its potential for ocean research.
@@ 3 OceanGPT
To obtain OceanGPT, we firstly construct the training corpus for ocean science and pre-train an ocean LLM based on LLaMA-2 (DBLP:journals/corr/abs-2307-09288 in Section 3.1. Then we propose DoInstruct, an automated framework for domain instruction generation to build an ocean domain-specific instruction dataset. Our framework leverages multi-agent collaboration and utilizes ocean literature to automatically generate a large volume of domain-specific instructions for ocean science tasks (Section 3.2). The overview training procedure of our OceanGPT is shown in Figure 2.
Figure 2: Overall training procedure of our OceanGPT.
@@ 3.1 Pre-training Stage
To pre-train the foundation model for ocean science tasks, it is essential to construct the pre-training corpus specific to ocean science. Therefore, we firstly collect a raw corpus of 67,633 documents from open-access literature. Specifically, we utilize the Python package pdfminer to convert the content of literature files into plain text. To ensure the quality and consistency of the data, further processing of the collected dataset is necessary. We apply regular expressions to filter out figures, tables, headers, footers, page numbers, URLs and references. Additionally, any extra spaces, line breaks, and other non-text characters are removed. The processed documents cover various aspects of ocean science such as ocean physics, ocean chemistry, ocean biology, geology, hydrology, etc. It is important to note that special characters, emoticons, and garbled characters are also replaced or eliminated during this process. We also employ hash-based methods to de-duplicate the data, which helps reduce the risk of over-fitting during pre-training and enhances its generalization capability. The detailed statistics is shown in Table LABEL:table:pre-training.
@@ 3.2 Domain Instruction Data Generation
Figure 3: Procedure of our proposed DoInstruct. We use agents as experts for each ocean topic and make them rapidly expand the instructions by collaboration. We design three agent roles: evolving generator, fine-tuned literature extractor and inspector with rule constraints.
@@
  • Life, culture and others. The ocean is not only a natural resource or a subject of scientific research; it is also an integral part of culture and lifestyle. This category consists of aspects ranging from history and culture to the mutual influences between the ocean and human societal activities, such as tourism, leisure, art, faith, and more.
@@
While these five topics are distinct, there might be some overlap as well. For instance, some issues related to ocean environmental protection might also be associated with the technology of ocean engineering. For the sake of convenience in data analysis, in the actual construction of the dataset, we map each sample to the most relevant category it belongs to.
@@
Agents as Domain (Ocean) Experts As shown in Figure 3 and Algorithm 1, we leverage agents as domain experts for each ocean topic and make them rapidly expand the instructions by collaboration. We first collect the seed instruction data and propose three strategies by using multiple agents acting as experts. The seed instruction dataset comes from annotations by ocean science researchers, which includes 5 major categories, over 500 sub-categories, and a total of more than 10,000 data entries.
@@
  • Evolving Agent as the Generator. We design an evolving approach that selects samples from the seed instruction dataset and simultaneously calls upon two agents (gpt-3.5-turbo) to evolve the selected samples. The evolution procedure includes two aspects: (1) we enrich the content of the sample by having the agent automatically add relevant background knowledge to it; (2) we guide the agent to refine the sample by conducting a more in-depth analysis of specific concepts or entities. Through multiple rounds of iterative execution, our method can rapidly expand the existing seed dataset, which allows for the rapid expansion of both the breadth and depth of information.
@@
  • Fine-Tuned Agent as the Literature Extractor. As shown in Figure 3, we collect a smaller expert-annotated corpus and use the BM25 to retrieve high quality sentences in a larger ocean corpus. We use the retrieved texts as high-quality candidate samples. Meanwhile, we fine-tune gpt-3.5-turbo with the seed instruction dataset, regarding the fine-tuned agent as the literature extractor. In other words, it can automatically extract questions (instruction) from the unsupervised ocean science corpus (output). Therefore, we utilize the agent to automatically build pairs of (instruction, output) on external ocean science literature.
@@
  • Agent as the Inspector with Rule Constraints. For the massively generated instructions, we use the pre-defined rules as constraints and perform filtering on the data. These rules include syntactic and semantic constraints as well as some basic definitions in the ocean domain. We describe these rules using natural language because many constraints and norms related to ocean science cannot be directly represented with expressions. Therefore, we provide prompts to the gpt-3.5-turbo API as demonstrations, letting it play the role of an Inspector. Our approach ensures that the generated ocean instruction data is of higher quality. The detailed prompt agent is shown in Table 3.
Figure 4: Statistics of our final instruction dataset. We use DoInstruct to expand more than 15,000 instructions (data-evolving, data-extracting) from the seed data (data-seed).
@@
Finally, we assign two extra gpt-3.5-turbo agents as roles to debate the quality of data and obtain high-quality instruction dataset. The designed framework is capable of rapidly constructing a ocean science dataset using multi-agents, and by incorporating external knowledge, it overcomes the limitations inherent to the agents themselves. Our framework can also be effectively applied to the instruction data construction in other scientific domains. It should be noted that we separately synthesize robot instructions to equip OceanGPT with the capability to interact with the environment. The statistics of our dataset are shown in Figure 4.
@@
We ask domain experts to carefully review and check data to ensure quality. Specifically, the human volunteers are first trained to make sure they have a comprehensive understanding of the task. Then we develop a platform that can help experts to sample instances from the instruction dataset. Next, the trained domain experts are asked to validate if there are potential errors in the sampled instances. The final IAA (inter-annotator agreement) score for our dataset is 0.82, which satisfies the research purpose.
@@ 4 Benchmarking Ocean Science Tasks
We provide detailed explanations of the experimental setup and the baseline models in Section 4.1. In Section 4.2, we construct an ocean-specific benchmark OceanBench to evaluate the capabilities of our OceanGPT. We describe the automatic and human evaluation in Section 4.2.
@@ 4.1 Implementation Details and Baselines
For the pre-training stage, we pre-train our OceanGPT based on the LLaMA-2 (DBLP:journals/corr/abs-2307-09288) for seven days with six A800 Nvidia GPUs. For the instruction-tuning stage, we employ the LoRA method (hu2021lora) to fine-tune the pre-trained model and choose three baseline models for comparison. We use the chat version of LLaMA-2 (Llama-2-7b-chat-hf [^2] https://huggingface.co/meta-llama/Llama-2-7b-chat-hf) , which is a generative language model optimized for dialogue use cases. We also use Vicuna-1.5333https://huggingface.co/lmsys/vicuna-7b-v1.5 (vicuna2023), a chat model which fine-tunes LLaMA-2 on dataset collected from ShareGPT. We further use ChatGLM2-6B [^4] https://huggingface.co/THUDM/chatglm2-6b, the optimized version of GLM (DBLP:conf/iclr/ZengLDWL0YXZXTM23). The detailed experimental settings are shown in Table 2 (Appendix A.3).
4.2 Datasets and Metrics
@@
To evaluate the capabilities of LLMs for oceanography tasks, we design a benchmark called OceanBench. Our benchmark includes a total of 15 type of ocean-related tasks, such as question-answering, extraction, and description tasks. Our evaluation samples are automatically generated from the seed dataset and have undergone deduplication and manual verification by domain experts. For the quality control, we further sample part of data and ask domain experts to evaluate the quality (those disagreed cases or bad cases will be manually fixed by domain experts.). The distribution of our OceanBench and the statistics can be found in Table 1.
@@
To evaluate the performance and reduce reliance on manual evaluation, we leverage GPT-4 as the evaluator for the experimental results. Inspired by wang2023mint; wang2023large, we utilize an effective calibration method to evaluate the performance of two LLMs. For each testing question, given two responses and from two LLMs, we query the GPT4 to obtain the comparison result. We note that LLMs are sensitive to the position of responses, therefore, alleviating the positional bias is very important. To balance the position bias, we exchange the order of the responses to form the new prompt. The final evaluating result is the sum of the test results for the two prompts with their order swapped.
@@
5 Evaluation

### v7

@@ Abstract
Ocean science, which delves into the oceans that are reservoirs of life and biodiversity, is of great significance given that oceans cover over 70% of our planet’s surface. Recently, advances in Large Language Models (LLMs) have transformed the paradigm in science. Despite the success in other domains, current LLMs often fall short in catering to the needs of domain experts like oceanographers, and the potential of LLMs for ocean science is under-explored. The intrinsic reasons are the immense and intricate nature of ocean data as well as the necessity for higher granularity and richness in knowledge. To alleviate these issues, we introduce OceanGPT, the first-ever large language model in the ocean domain, which is expert in various ocean science tasks. We also propose DoInstruct, a novel framework to automatically obtain a large volume of ocean domain instruction data, which generates instructions based on multi-agent collaboration. Additionally, we construct the first oceanography benchmark, OceanBench, to evaluate the capabilities of LLMs in the ocean domain. Though comprehensive experiments, OceanGPT not only shows a higher level of knowledge expertise for oceans science tasks but also gains preliminary embodied intelligence capabilities in ocean technology.
OceanGPT: A Large Language Model for Ocean Science Tasks
Zhen Bi1,2,5,6,  Ningyu Zhang1,2,511footnotemark: 1,  Yida Xue1,  Yixin Ou1, Daxiong Ji2,3Guozhou Zheng2,4,  Huajun Chen1,2thanks: Corresponding author. 1College Computer Science and Technology, Zhejiang University  2Donghai Laboratory 3Ocean College, Zhejiang University 4Zhoushan-Zhejiang University Ocean Research Center 5School of Software Technology, Zhejiang University, 6Huzhou University {bizhen_zju,zhangningyu,ouyixin,jidaxiong,guozhou,huajunsir}@zju.edu.cn Project Website: https://oceangpt.zjukg.cn/
@@ 1 Introduction
Figure 1: Capabilities of OceanGPT. Our proposed model not only shows a higher level of knowledge expertise for oceans science tasks but also gains preliminary embodied intelligence capabilities in ocean technology.
Ocean science, which delves into the intricacies of oceans that cover over 70% of our planet’s surface, is essential not only for understanding the rich reservoirs of life and biodiversity but also for recognizing their pivotal role in regulating the global climate and supporting economies (Esaias et al., 1998; Falkowski, 2012; Visbeck, 2018; Jin et al., 2023). Recently, advances in Large Language Models (LLMs) (OpenAI, 2023; Jiang et al., 2023; Zha et al., 2023; Yin et al., 2023; Zhao et al., 2023) have transformed the paradigm in science domains such as medical science (Moor et al., 2023), molecular science (Fang et al., 2023), protein science (Lin et al., 2023) and geoscience (Deng et al., 2023). However, the potential for the large language model in ocean science is under-explored.
@@
Through extensive experiments, OceanGPT shows superiority for diverse ocean science tasks. Note that our benchmark data is based on criteria manually evaluated by ocean experts, and can accurately reflect the capabilities that LLMs possess in the field of ocean science. As depicted in Figure 1, our model can comprehensively answer questions according to the instructions of oceanographers, which demonstrates its expertise in oceanography. We further explore the potential of OceanGPT from the perspectives of ocean engineering. Specifically, we integrate ocean robotics instructions into the training data and evaluate its ability via code or console commands. OceanGPT not only demonstrates a higher level of knowledge expertise but also gains preliminary embodied intelligence capabilities in ocean technology.
@@ 2 Related Work
The landscape of LLM (Brown et al., 2020; Chowdhery et al., 2022; Touvron et al., 2023a, b) has rapidly evolved and achieved a series breakthroughs. Rae et al. (2021); Zhang et al. (2022); Thoppilan et al. (2022); Scao et al. (2022); Zeng et al. (2023) have explored the performance across a wide range of model scales and broadened the application scope. (Qiao et al., 2023a; Zhang et al., 2023a; Qiao et al., 2023b; Wang et al., 2023a; Xi et al., 2023). Retrieval-Augmented Generation (RAG) is a useful solution by incorporating knowledge from external databases (Gao et al., 2023; Lewis et al., 2020; Schick et al., 2023; Khandelwal et al., 2020). To align LLMs, instruction tuning (Wei et al., 2022; Zhang et al., 2023b; Ouyang et al., 2022; Taori et al., 2023; Wang et al., 2023d; Chiang et al., 2023; Xu et al., 2023) is a crucial technique to alignment with user preferences and desired outputs. Different from those, we train a totally new ocean science large language model and introduce an effective domain instruction generation framework via multi-agent collaboration.
@@
LLMs have emerged as cornerstone models in addressing challenges within scientific research. Singhal et al. (2022) explores the potential of clinical LLMs and introduces a human evaluation framework and instruction prompt tuning. Moor et al. (2023) proposes generalist medical AI that is capable of handling diverse medical tasks using self-supervised learning on large datasets. Kraljevic et al. (2021) introduces MedGPT, a model using EHR data and Named Entity Recognition tools for predicting future medical events. BioGPT (Luo et al., 2022) is a language model pre-trained on biomedical literature for improved text generation and mining. Theodoris et al. (2023) describes Geneformer, a model pre-trained on single-cell transcriptomes for making predictions with limited data in network biology. Lin et al. (2023) demonstrates the prediction of atomic-level protein structure from primary sequences using scaled-up language models. Deng et al. (2023) introduces the first LLM specifically designed for geoscience, including its training and benchmarking protocols. Chen et al. (2023) presents tele-knowledge pre-training for fault analysis. Different from previous works, we design the first large language model for ocean science tasks and explore its potential for ocean research.
@@ 3 OceanGPT
To obtain OceanGPT, we firstly construct the training corpus for ocean science and pre-train an ocean LLM based on LLaMA-2 Touvron et al. (2023b) in Section 3.1. Then we propose DoInstruct, an automated framework for domain instruction generation to build an ocean domain-specific instruction dataset. Our framework leverages multi-agent collaboration and utilizes ocean literature to automatically generate a large volume of domain-specific instructions for ocean science tasks (Section 3.2). The overview training procedure of our OceanGPT is shown in Figure 2.
@@ 3.1 Pre-training Stage
Figure 2: Overall framework of OceanGPT.
To pre-train the foundation model for ocean science tasks, it is essential to construct the pre-training corpus specific to ocean science. Therefore, we firstly collect a raw corpus of 67,633 documents from open-access literature. For the specific volumes we choose, we prefer to consider publications from recent years to ensure the inclusion of the latest research and developments. At the same time, we select some historically significant literature to help the LLM understand the developmental history of the field. For diversity, we choose articles from different sources to ensure coverage of various research perspectives and methods. Specifically, we utilize the Python package pdfminer to convert the content of literature files into plain text. To ensure the quality and consistency of the data, further processing of the collected dataset is necessary. We apply regular expressions to filter out figures, tables, headers, footers, page numbers, URLs and references. Additionally, any extra spaces, line breaks, and other non-text characters are removed. The processed documents cover various aspects of ocean science such as ocean physics, ocean chemistry, ocean biology, geology, hydrology, etc. It is important to note that special characters, emoticons, and garbled characters are also replaced or eliminated during this process. We also employ hash-based methods to de-duplicate the data, which helps reduce the risk of over-fitting during pre-training and enhances its generalization capability.
@@ 3.2 Domain Instruction Data Generation
Figure 3: Procedure of our proposed DoInstruct. We use agents (gpt-3.5-turbo) as experts for each ocean topic and make them rapidly expand the instructions by collaboration. In this framework, we design three agent roles: evolving generator, fine-tuned literature extractor and inspector with rule constraints.
@@
  • Life, culture and others. The ocean is not only a natural resource or a subject of scientific research; it is also an integral part of culture and lifestyle. This category consists of aspects ranging from history and culture to the mutual influences between the ocean and human societal activities, such as tourism, leisure.
@@
While these five topics are distinct, there might be some overlap as well. For instance, some issues related to ocean environmental protection might also be associated with the technology of ocean engineering. For the sake of convenience in data analysis, in the actual construction of the dataset, we map each sample to the most relevant category.
@@
Agents as Domain (Ocean) Experts. In Figure 3, we use agents as domain experts for each ocean topic and make them rapidly expand the instructions by collaboration. We collect the seed instruction data and propose three strategies by using multiple agents acting as experts.
To construct the seed dataset, we employ dozens of annotators with rich backgrounds in marine science. Each annotator is responsible for several topics and they first manually write some representative example for each marine topic. Then we use LLMs to mimic the existing data to generate a large number of similar samples. All samples are ultimately manually checked by the annotators. The entire process is very time-consuming, with all the experts spending a total of four days to validate the seed data. The final seed instruction dataset includes 5 major categories, over 500 sub-categories and a total of more than 10,000 data samples.
Algorithm 1 Domain Instruction Data Generation
1:Seed dataset $S$ with format $({inst,output})$, Ocean literature corpus $O$, Pre-defined rules $R$ for filtering
2:High-quality instruction dataset $H$
3:Initialize empty datasets.
4:${Step1Data}=\emptyset,{Step2Data}=\emptyset,H=\emptyset$
5$\triangleright$ Agent Collaboration as Domain Experts.
6:for each sample in $S$ do
7:     ${inst,output}\leftarrow{sample}$
8:     ${enriched\_sample}\leftarrow{Enrich(inst,output)}$
9:     ${refined\_sample}\leftarrow{Refine(inst,output)}$
10:     ${Step1Data}\leftarrow{Step1Data}\cup{enriched\_sample}\cup{refined\_sample}$
11:end for
12$\triangleright$ Fine-Tuned Agent as Literature Extractor.
13:${RetrievedTexts}\leftarrow{BM25\_Retrieve(O)}$
14:${Model}~{}M\leftarrow{FineTune(S_{reverse})}$
15:for each document in ${RetrievedTexts}$ do
16:     ${output}\leftarrow{document.content}$
17:     ${inst}\leftarrow{M(output)}$
18:     ${Step2Data}\leftarrow{Step2Data}\cup({inst,output})$
19:end for
20$\triangleright$ Agent as Inspector with Rule Constraints.
21:${MergedData}\leftarrow{Inspector(Step1Data,Step2Data,R)}$
22$\triangleright$ Quality Control by Debating.
23:for each sample in MergedData do
24:     ${inst,output}\leftarrow{sample}$
25:     ${debate\_result}\leftarrow{Debate(inst,output)}$
26:     if debate_result is high-quality then
27:         $H\leftarrow H\cup{sample}$
28:     end if
29:end for
30:return $H$
@@
  • Evolving Agent as the Generator. We design an evolving approach that selects samples from the seed dataset and simultaneously calls upon two agents (gpt-3.5-turbo) to evolve the selected samples. The evolution procedure includes two aspects: (1) we enrich the content of the sample by having the agent automatically add relevant background knowledge to it; (2) we guide the agent to refine the sample by conducting a more in-depth analysis of specific concepts or entities. Through multiple rounds of iterative execution, our method can rapidly expand the existing seed dataset, which allows for the rapid expansion of both the breadth and depth of information.
@@
  • Fine-Tuned Agent as the Literature Extractor. As shown in Figure 3, we collect a smaller expert-annotated corpus and use the BM25 to retrieve high quality sentences in a larger ocean corpus. We regard the retrieved texts as high-quality candidate samples. Meanwhile, we fine-tune gpt-3.5-turbo with the seed instruction dataset, regarding the fine-tuned agent as the literature extractor. In other words, it can automatically extract instructions (inst) from the unannotated ocean science corpus (output). Therefore, we utilize the agent to automatically build pairs of (inst, output) on external ocean science literature.
@@
  • Agent as the Inspector with Rule Constraints. For the massively generated instructions, we use the pre-defined rules as constraints and perform filtering on the data. These rules include syntactic and semantic constraints as well as basic definitions in the ocean domain. We describe these rules using natural language because many constraints and norms related to ocean science cannot be directly represented with expressions. Therefore, we provide prompts to the gpt-3.5-turbo API as demonstrations, letting it play the role of an inspector. Our method ensures that the generated ocean instruction data is of higher quality. Detailed prompt is shown in Table 5.
@@
Finally, we assign two extra gpt-3.5-turbo agents as roles to debate the quality of data and obtain high-quality instruction dataset. Our designed framework can rapidly constructing a ocean science dataset using multi-agents, and by incorporating external knowledge from marine literature, it overcomes the limitations inherent to the agents themselves. Our framework can also be effectively applied to the instruction data construction in other scientific domains. It should be noted that we separately synthesize robot instructions to equip OceanGPT with the capability to interact with the environment. The procedure is in Algorithm 1 and the statistics of dataset is in Figure 4.
Figure 4: Statistics of our final instruction dataset. We use DoInstruct to expand more than 150,000 instructions (data-evolving, data-extracting).
@@
We ask domain experts to carefully review and check data to ensure quality. Specifically, the human volunteers are first trained to make sure they have a comprehensive understanding of the task. Then, we develop a platform that can help experts to randomly sample 10% instances from the generated instruction dataset. Next, the trained domain experts are asked to validate if there are potential errors in the sampled instances. The final IAA (inter-annotator agreement) score for our dataset is 0.82, which satisfies the research purpose.
@@ 4 Benchmarking Ocean Science Tasks
We provide detailed explanations of the experimental setup and the baseline models in Section 4.1. In Section 4.1, we construct an ocean-specific benchmark OceanBench [^1] https://huggingface.co/datasets/zjunlp/OceanBench to evaluate the capabilities of our OceanGPT. We describe the automatic and human evaluation in Section 1.
@@ 4.1 Implementation Details and Baselines
For the pre-training stage, we pre-train our OceanGPT [^2] We release the 7B version at https://huggingface.co/zjunlp/OceanGPT-7B-v0.1 for evaluation, and further release the 2B (MiniCPM-2B-sft-bf16) and 8B (Meta-Llama-3-8B-Instruct) version at https://huggingface.co/collections/zjunlp/oceangpt-664cc106358fdd9f09aa5157. based on the LLaMA-2 (Touvron et al., 2023b) for seven days with six A800 Nvidia GPUs. For the instruction-tuning stage, we employ the LoRA method (Hu et al., 2021) to fine-tune the pre-trained model and choose three baseline models for comparison. We use the chat version of LLaMA-2 (Llama-2-7b-chat-hf) , which is a generative language model optimized for dialogue use cases. We also use Vicuna-1.5 (Chiang et al., 2023), a chat model which fine-tunes LLaMA-2 on dataset collected from ShareGPT. We further use ChatGLM2-6B, the optimized version of GLM (Zeng et al., 2023). The detailed experimental settings are shown in Table 2 (Appendix A).
@@
To evaluate the capabilities of LLMs for oceanography tasks, we design a benchmark called OceanBench. Our benchmark includes a total of 15 ocean-related tasks such as question-answering, extraction, and description. Our evaluation samples are automatically generated from the seed dataset and have undergone deduplication  [^3] We also perform deduplication between the benchmark and our training dataset to avoid the data leakage in the training stage of OceanGPT. The detailed explanation about the similarity calculating deduplication method is in Appendix The Similarity Calculating Method in the Deduplication Procedure. and manual verification by experts.
For the quality control, we further sample part of data and ask domain experts to evaluate the quality (those disagreed cases or bad cases will be manually fixed by domain experts.). The distribution of our designed OceanBench and the detailed statistics can be found in Table 1 and Figure 11.
Task Num Task Num
Analysis 674 Classification 895
Judgment 655 Letter Writing 359
Open-ended Generation 930 Extraction 1,078
Recommendation 1,089 Description 1,246
Summary 149 Editing 1,075
Identification 464 Transformation 401
Question Answering 1,230 Others 157
Commonsense Reasoning 1,024
Table 1: The detailed statistics of OceanBench.
Metrics.
For the task-level calculation, we compare the effectiveness of two models for each task. When one model performs better on the majority of test samples in a single task, it is considered to ’win’ that task. For the instance-level computation process, we do not differentiate between specific tasks and instead calculate overall metrics.
@@
To evaluate the performance and reduce reliance on manual evaluation, we leverage GPT-4 as the evaluator. Inspired by Wang et al. (2023c, b), we utilize an effective calibration method to evaluate the performance of two LLMs. For each testing question, we query the GPT4 to obtain the comparison result when given two outputs from two LLMs. We note that LLMs are sensitive to the position of responses, so alleviating the positional bias is very important. To balance the position bias, we exchange the order of the responses to form the new prompt. The final evaluating result is the sum of the test results for the two prompts with their order swapped.
@@
5 Results
Figure 5: Ocean task-level results. Left: Automatic evaluation. Right: Human evaluation. Compared to baselines, OceanGPT performs better than llama2-chat-7b, vicuna-1.5-7b and chatglm2-6b in both two settings. The instance-level result is in Figure 10 (Appendix A).